OpenStudy (anonymous):

hey guys, i got a question.... is energy detectable without mass?..... if yes.. wat is the meaning of energy in empty sace?

OpenStudy (anonymous):

hm from what i know, Kinetic Energy and Potential Energy have Mass in their formula KE = (1/2)MV^2 PE = Mgh E = MC^2